# Quantifying the dynamic structural resilience of international staple food trade networks: An entropy-based approach

### v2

@@ 2.3 Resilience of the international food trade system
Based on these two system properties, the ratio $\alpha$, a more comprehensive metric to indicate the order of a system, is proposed for reflecting the trade-off between efficiency and redundancy in the food trade system, which is expressed as
@@
where $0\leq\alpha\leq 1$. Inspired by ecological systems that their orders are all close to $1/{\mathrm{e}}$ (Zorach and Ulanowicz, 2003), Ulanowicz (2009) defined the fitness $R$ of a system for change to be the product of $\alpha$ and the Boltzmann measure of its disorder such that $R=-c\alpha\ln\alpha$, where $c$ is an appropriate scalar constant. When $\alpha=1/{\mathrm{e}}$, $R^{\prime}=0$, which means the order of the system is optimized. The underlying assumption is that ecosystems exhibit superior trade-off because they have undergone long-term natural selection (Ulanowicz, 2009; Liang et al., 2020; Kharrazi et al., 2013). Subsequently, let $c=1$, the resilience $R$ of a food trade network can be defined as
@@
where the maximum value is $1/{\mathrm{e}}\approx 0.3679$. Hence, a resilient food trade system possesses both the capacity to flexibly allocate food (efficient) and to absorb shocks via several pathways (redundant). When $\alpha>1/{\mathrm{e}}$, the system is more efficient and productive but more vulnerable, and vice versa it is more redundant but more inefficient. Notably, resilience vanishes when $\alpha=0$ (overly redundant) or $\alpha=1$ (overly efficient).
@@ 2.4 Relative contributions of changes in efficiency and redundancy
[…]
@@
Subsequently, we define a ratio as follows (Liang et al., 2020):
@@ 2.5 Relative contributions of changes in P, PMI, and PMR
The second internal component is the inter-dependency among economies, defined as a point-wise mutual information (PMI) matrix containing element $\mathrm{pmi}_{ij}$ that denotes the degree of inter-dependency between economies $i$ and $j$ (Kharrazi and Fath, 2016). A higher value indicates a higher probability of a flow between given economies. The element $\mathrm{pmi}_{ij}$ of PMI matrix is expressed as:
@@
[…]
@@
[…]
@@
2.6 Significance of economies and trade relationships in resilience
The perturbation methods, including node removal, link removal, and the global weight perturbation, are commonly employed to investigate changes in system resilience (Liu et al., 2024a). Since the global weight perturbation does not affect results in Eqs. (1) and (5), according to the leave-one-out approach (Hué et al., 2019; Zedda and Cannas, 2020), we quantify the contributions of changes in economies and relationships to changes in resilience to identify critical economies and trade relationships from the resilience perspective.
@@ 2.6 Significance of economies and trade relationships in resilience
Similarly, for a given trade relationship $e_{ij}$, we consider a system $G_{(-e_{ij})}$ with $N_{\mathcal{E}}-1$ links that includes all relationships except the relationship $e_{ij}$ and denote the resilience of this system as $R^{G_{(-e_{ij})}}$. Hence, the contributions of changes in relationship $e_{ij}$ can be expressed as
$\Delta R_{e_{ij}}=\frac{\lvert R^{G}-R^{G_{(-e_{ij})}}\rvert}{R^{G}},$ (29)
which also denotes the significance of relationship $e_{ij}$ in the international food trade system from a network resilience perspective. A higher $\Delta R_{e_{ij}}$ implies that relationship $e_{ij}$ plays a more important role in system resilience.
@@ 2.7 Resilience of core and peripheral food trade networks
The international food trade network exhibits a pronounced core-periphery structure (Dong et al., 2018), which prompts us to investigate the differences between resilience of core and peripheral food trade networks. It should be emphasized that our study explicitly excludes cumulative economy removal scenarios, as termination of trade relationships remains operationally feasible (Guo et al., 2025), which fundamentally differs from complete embargoes on economy’s food trade. We sort all trade relationships either in ascending or descending order of $\Delta R_{e_{ij}}$ and define $p=\frac{m}{N_{\mathcal{E}}}$ as the ratio of the number of relationships within the ordered set $\mathcal{E}^{\prime}$ to the total number of relationships $N_{\mathcal{E}}$, where the set of given trade relationships =$\mathcal{E}^{\prime}=\{e_{1},e_{2},\dots,e_{m}\}$ and $1\leq m\leq N_{\mathcal{E}}$.
Hence, $G_{(-p)}$ is defined as the food trade network that includes the proportion $1-p$ of relationships in order of $\Delta R_{e_{ij}}$. For ascending order of $\Delta R_{e_{ij}}$, the network $G_{(-p)}$ can be regarded as a core food trade network when the value of $p$ is large (that is, the network is constituted by relationships with greater impact on resilience). Also, for descending order of $\Delta R_{e_{ij}}$, the network $G_{(-p)}$ can be regarded as a peripheral food trade network when the value of $p$ is large (that is, the network is constituted by relationships with less impact on resilience). Subsequently, we can calculate the resilience of core or peripheral food trade networks $R^{G_{(-p)}}$ with respect to the changes in $p$, which may help clarify the role of the above networks in sustaining network resilience and guide structure optimization for the international food trade.
@@ 3 Dynamic evolution of the international food trade network and its resilience
Figure 1: The evolution of topological properties of international food trade networks. a,b, The numbers of nodes $N_{\mathcal{V}}$ and links $N_{\mathcal{E}}$ are two alternative measures of network size. c,d, Average degree $\langle k\rangle$ and average strength $\langle s\rangle$ represent the average number of trade partners and trade volume of economies, respectively. e, Average clustering coefficient $\langle C\rangle$ represents the average actual connections among the neighbors of individual nodes. f, Density $\rho$ reflects the level of interconnectedness among economies. g, Average shortest path length $L$ represents the level of separation among economies. h, Diameter $D$ represents the maximum of shortest path lengths between any two economies. i, Modularity $Q$ represents the denseness of flows within a community compared to between communities.
Given that maize, rice, soybean, and wheat are foods that account for most of the calories consumed by the world’s population (D’Odorico et al., 2014; Zhang and Zhou, 2023), we select the international bilateral trade quantities of maize (CPC code [^1] Central Product Classification (CPC) code, a product classification for goods and services promulgated by the United Nations Statistics Division.: 0112), rice (CPC code: F0030), soybean (CPC code: 0141), and wheat (CPC code: 0111) from 1986 to 2022, provided by the “detailed trade matrix” dataset [^2] FAOSTAT, www.fao.org/faostat/en/#data/TM. of the FAO to construct networks, where economies are considered as nodes and trade relationships between economies serve as directed and weighted links with the trade quantities being the weights of links. We note that the abbreviations for economies used in the figures and tables are taken from the International Standard Organization [^3] ISO, www.iso.org/iso-3166-country-codes.html..
@@
Specifically, we analyze the international trade networks of four foods in 2022, as shown in Fig. 2. The node size refers to its total trade volume (including imports and exports) and we label the top 10 economies by the size (Fig. 2a-d). Compared to maize and rice, the soybean and wheat trade networks exhibit sparser connectivity and lower density (Fig. 1f), manifested by comparable numbers of participating economies but fewer trade relationships (Fig. 1a,b). For both of them, the first quartile is zero in the logarithmic-transformed boxplots of $k^{\mathrm{out}}$ and $s^{\mathrm{out}}$ (Fig. 2f,h), indicating their more concentrated exporting characteristics. The difference, however, is that trade volume of wheat varies less among importers and is more evenly distributed (Fig. 2g). Moreover, the soybean trade network exhibits substantial heterogeneity in the size of participating economies, indicating highly monopolistic characteristics by a few economies such as the United States, Brazil, and China (Fig. 2c).
Overall, evidently, the United States and China are the most important trading economies, with the former being a major exporter and the latter a major importer. With regard to exports, the United States has vast arable land and suitable climatic condition for the large-scale cultivation of a wide range of foods. Not only in production but also in variety, the adequate level of supply has also contributed to it being a major exporter. With regard to imports, although China has a large area of arable land, it has a large population and a small per capita area of arable land, which causes China having to rely on imports to meet domestic demands.
@@
Furthermore, we calculate efficiency, redundancy, $\alpha$ values, and resilience of the international food trade networks from 1986 to 2022 (Fig. 3a-d). Overall, trade networks of maize, rice, and wheat are more resilient. Separately, maize and wheat exhibit parallel enhancements in both efficiency and redundancy, resulting in the better trade-off within $\alpha$ framework. Rice exhibits enhancement in redundancy but fluctuation in efficiency, resulting in a slowly declining resilience in recent years. Conversely, the soybean trade network is not only more vulnerable, its resilience fluctuates more significantly. As is known, approximately 87% of soybean export quantities originated from Brazil and the United States in 2022. Due to the excessively high concentration of exporters, most economies, except for the main exporters, remain relatively isolated, with limited alternative trade pathways. Consequently, the soybean trade network lacks both efficiency and redundancy, and exhibits an insufficient trade-off within the $\alpha$ framework, ultimately resulting in lower resilience. Notably, $\alpha$ values of all networks are less than $1/{\mathrm{e}}$, resulting in the highly positive correlation between efficiency and resilience (Fig. 3e) and the negative correlation between redundancy and resilience (Fig. 3f). It indicates the current trade networks possess sufficient redundancy but require enhanced efficiency, especially for soybean.
4 Network properties and resilience of the international food trade network
4.1 Correlations between topological properties and resilience
We investigate the correlations between topological properties of the international food trade networks and their resilience, as shown in Fig. 4. Regardless of foods, evidently, we find that modularity exhibits the strongest positive correlation to resilience, indicating that more resilient states of international trade may be associated with stronger community structures (Garlaschelli and Loffredo, 2005; Li et al., 2024). A higher modularity reflects denser intra-community flows relative to inter-community flows, corresponding to processes of regional integration that leads to stronger inter-dependence among economies within the community and enhanced overall network efficiency (de Raymond et al., 2021). For maize and wheat, for example, the correlation coefficients between resilience and modularity are 0.72 and 0.80, respectively.
@@ 4.1 Correlations between topological properties and resilience
Figure 4: The correlations between topological properties of the international food trade networks and their resilience. a-i, Scattered distribution of resilience and topological indicators. In this figure, the red dotted line indicates $1/{\mathrm{e}}$. j, Kendall’s correlation coefficients between resilience and topological indicators. In this figure, $R_{m}$, $R_{r}$, $R_{s}$, and $R_{w}$ denote resilience of trade networks of maize, rice, soybean, and wheat, respectively.
The correlation between resilience and other topological properties varies across different foods. For example, the network size and clustering characteristics are positively correlated with resilience for maize, soybean, and wheat, but are negatively correlated with resilience for rice. In mature trade networks such as those for maize and wheat, evident clustering characteristics enhance trade efficiency (e.g., through regional alliances and international cooperation organizations, such as NAFTA [^6] NAFTA, The North American Free Trade Area. and MERCOSUR [^7] MERCOSUR, The Southern Common Market.), while the larger network size increases systemic redundancy. Together, these features jointly contribute to sustaining network resilience. This finding fully supports the view that internationalization does not undermine, but rather strengthens the resilience of mature food trade networks (Sartori and Schiavo, 2015).
@@
However, resilience of the rice trade network appears to decline with increases in both clustering coefficients and the network size, potentially due to its inherent network characteristics and the nature of rice as a commodity. The network is heavily concentrated around a few major exporters (such as India, Thailand, and Vietnam) and importers (such as China and the Philippines), forming a single-core structure of Asia. Despite greater connectivity, this centralized architecture has exacerbated systemic vulnerability during network expansion. Meanwhile, while geographic proximity suggests potential trade alliances, the lack of strategic complementarity, exemplified by competitive underpricing between Vietnam and Thailand, undermines cluster resilience by fostering internal fragmentation. Additionally, as an inelastic staple commodity with low substitution elasticity, the rice trade network faces catastrophic systemic collapse once major Asian producers implement export restrictions to ensure domestic food security (e.g., India’s 2022 export ban for rice).

### v3

@@ 2.3 Resilience of the international food trade system
It is well established that improving resilience requires a balance between efficiency (or flexibility) and redundancy (or diversity) (Zhu et al., 2025; Ataburo et al., 2024; van Staveren, 2023; Essuman et al., 2020; Brede and de Vries, 2009; Kamalahmadi et al., 2022). Based on these two system properties, the ratio $\alpha$, a more comprehensive metric to indicate the order of a system, is proposed for reflecting the trade-off between efficiency and redundancy in the food trade system, which is expressed as
@@
where $0\leq\alpha\leq 1$. Inspired by ecological systems that their orders are all close to $1/{\mathrm{e}}$ (Zorach and Ulanowicz, 2003), Ulanowicz (2009) defined the fitness $F$ of a system for change to be the product of $\alpha$ and the Boltzmann measure of its disorder such that $F=-c\alpha\ln\alpha$, where $c$ is an appropriate scalar constant. When $\alpha=1/{\mathrm{e}}$, $F^{\prime}=0$, which means the order of the system is optimized. The underlying assumption is that ecosystems exhibit superior trade-off because they have undergone long-term natural selection (Ulanowicz, 2009; Liang et al., 2020; Kharrazi et al., 2013). Subsequently, let $c=1$, the resilience $R$ of a food trade network can be defined as
@@
where the maximum value is $1/{\mathrm{e}}\approx 0.3679$. Hence, a resilient food trade system possesses both the capacity to flexibly allocate food (efficient) and to absorb shocks via several pathways (redundant). When $\alpha>1/{\mathrm{e}}$, the system is more efficient and productive but more vulnerable, and vice versa it is more redundant but more inefficient. Notably, resilience vanishes when $\alpha=0$ (overly redundant) or $\alpha=1$ (overly efficient), and more details are provided in Appendix A.
@@ 2.4 Relative contributions of changes in efficiency and redundancy
$\begin{aligned} \Delta R&=\frac{1+\ln(e)-\ln(e+r)}{\left(e+r\right)^{2}}(e\Delta r-r\Delta e)\\ &=\left(\frac{1+\ln(e)-\ln(e+r)}{\left(e+r\right)^{2}}e\Delta r\right)-\left(\frac{1+\ln(e)-\ln(e+r)}{\left(e+r\right)^{2}}r\Delta e\right)\\ &\triangleq\Delta_{r}R+\Delta_{e}R\end{aligned},$ (10)
@@
As only discrete data are available in empirical studies, we approximate $e$ and $r$ by the arithmetic mean of their respective values for base year and observation year (Ang and Zhang, 2000). Subsequently, we can define a ratio as follows (Liang et al., 2020):
@@ 2.5 Relative contributions of changes in P, PMI, and PMR
The second internal component is the inter-dependency among economies, defined as a point-wise mutual information (PMI) matrix containing element $\mathrm{pmi}_{ij}$ that denotes the degree of inter-dependency between economies $i$ and $j$ (Kharrazi and Fath, 2016). A higher value indicates a higher probability of a flow between given economies. The element $\mathrm{pmi}_{ij}$ of PMI matrix is expressed as
@@
$\begin{aligned} \Delta e_{ij}&=\Delta p_{ij}\times\mathrm{pmi}_{ij}+p_{ij}\times\Delta\mathrm{pmi}_{ij}\\ &=\frac{1}{2}\left(\Delta p_{ij}\times\mathrm{pmi}_{ij}^{0}+\Delta p_{ij}\times\mathrm{pmi}_{ij}^{t}\right)+\frac{1}{2}\left(p_{ij}^{0}\times\Delta\mathrm{pmi}_{ij}+p_{ij}^{t}\times\Delta\mathrm{pmi}_{ij}\right)\\ &=\Delta p_{ij}\times\frac{1}{2}\left(\mathrm{pmi}_{ij}^{0}+\mathrm{pmi}_{ij}^{t}\right)+\Delta\mathrm{pmi}_{ij}\times\frac{1}{2}\left(p_{ij}^{0}+p_{ij}^{t}\right)\end{aligned},$ (15)
@@
$\begin{aligned} \Delta R&=\Delta_{e}R+\Delta_{r}R\\ &=\mathrm{peR}\times\Delta e+\mathrm{prR}\times\Delta r\\ &=(\mathrm{peR}\times\Delta_{p}e+\mathrm{prR}\times\Delta_{p}r)+(\mathrm{peR}\times\Delta_{\mathrm{pmi}}e)+(\mathrm{prR}\times\Delta_{\mathrm{pmr}}r)\\ &\triangleq\Delta_{p}R+\Delta_{\mathrm{pmi}}R+\Delta_{\mathrm{pmr}}R\end{aligned},$ (27)
@@
2.6 Relative contributions of changes in economies and trade relationships
The perturbation methods, including node removal, link removal, and the global weight perturbation, are commonly employed to investigate changes in system resilience (Liu et al., 2024a). Since the global weight perturbation does not affect results in Eqs. (1) and (5), according to the leave-one-out approach (Hué et al., 2019; Zedda and Cannas, 2020), we employ node removal and link removal to quantify the contributions of changes in economies and relationships to changes in resilience, thereby identifying critical economies and trade relationships from the resilience perspective.
@@ 2.6 Relative contributions of changes in economies and trade relationships
Similarly, for a given trade relationship $\varepsilon_{ij}$, we consider a system $G_{(-\varepsilon_{ij})}$ with $N_{\mathcal{E}}-1$ links that includes all relationships except the relationship $\varepsilon_{ij}$ and denote the resilience of this system as $R^{G_{(-\varepsilon_{ij})}}$. Hence, the contributions of changes in relationship $\varepsilon_{ij}$ can be expressed as
$\Delta R_{\varepsilon_{ij}}=\frac{\lvert R^{G}-R^{G_{(-\varepsilon_{ij})}}\rvert}{R^{G}},$ (29)
which also denotes the significance of relationship $\varepsilon_{ij}$ in the international food trade system from a network resilience perspective. A higher $\Delta R_{\varepsilon_{ij}}$ implies that relationship $\varepsilon_{ij}$ plays a more important role in system resilience.
@@ 2.7 Resilience of core and peripheral food trade networks
The international food trade network exhibits a pronounced core-periphery structure (Dong et al., 2018), which prompts us to investigate the differences between resilience of core and peripheral food trade networks. It should be emphasized that our study explicitly excludes cumulative economy removal scenarios, as termination of trade relationships remains operationally feasible (Guo et al., 2025), which fundamentally differs from complete embargoes on economy’s food trade. We sort all trade relationships either in ascending or descending order of $\Delta R_{\varepsilon_{ij}}$. In our analysis, a proportion $p$ of the relationships are first removed from the ordered set and $G_{(-p)}$ is defined as the food trade network that includes the remaining proportion 1p1-p of relationships in order of $\Delta R_{\varepsilon_{ij}}$. For ascending order of $\Delta R_{\varepsilon_{ij}}$, the network $G_{(-p)}$ can be regarded as a core food trade network when the value of $p$ is large (that is, the network is constituted by relationships with greater impact on resilience). Also, for descending order of $\Delta R_{\varepsilon_{ij}}$, the network $G_{(-p)}$ can be regarded as a peripheral food trade network when the value of $p$ is large (that is, the network is constituted by relationships with less impact on resilience). Subsequently, we can calculate the resilience of core or peripheral food trade networks $R^{G_{(-p)}}$ with respect to the changes in $p$, which may help clarify the role of the above networks in sustaining network resilience and guide structure optimization for the international food trade.
@@ 3 Dynamic evolution of the international food trade network and its resilience
Given that maize, rice, soybean, and wheat are foods that account for most of the calories consumed by the world’s population (D’Odorico et al., 2014; Zhang and Zhou, 2023), we select the international bilateral trade quantities of maize (CPC code [^1] Central Product Classification (CPC) code, a product classification for goods and services promulgated by the United Nations Statistics Division.: 0112), rice (CPC code: F0030), soybean (CPC code: 0141), and wheat (CPC code: 0111) from 1986 to 2022, provided by the “detailed trade matrix” dataset [^2] FAOSTAT, www.fao.org/faostat/en/#data/TM. of the FAO, to construct networks, where economies are considered as nodes and trade relationships between economies serve as directed and weighted links with the trade quantities being the weights of links. We note that the abbreviations for economies used in the figures and tables are taken from the International Standard Organization [^3] ISO, www.iso.org/iso-3166-country-codes.html..
Figure 1: The evolution of topological indicators of international food trade networks. a,b, The numbers of nodes $N_{\mathcal{V}}$ and links $N_{\mathcal{E}}$ are two alternative measures of network size. c,d, Average degree $\langle k\rangle$ and average strength $\langle s\rangle$ represent the average number of trade partners and trade volume of economies, respectively. e, Average clustering coefficient $\langle C\rangle$ represents the average actual connections among the neighbors of individual nodes. f, Density $\rho$ reflects the level of interconnectedness among economies. g, Average shortest path length $L$ represents the level of separation among economies. h, Diameter $D$ represents the maximum of shortest path lengths between any two economies. i, Modularity $Q$ represents the denseness of flows within a community compared to between communities.
@@
Specifically, we analyze the international trade networks of four foods in 2022, as shown in Fig. 2. The node size refers to its total trade volume (including imports and exports) and we label the top 10 economies by the size (Fig. 2a-d). Evidently, the United States and China are the most important trading economies, with the former being a major exporter and the latter a major importer. With regard to exports, the United States has vast arable land and suitable climatic condition for the large-scale cultivation of a wide range of foods. Not only in production but also in variety, the adequate level of supply has also contributed to it being a major exporter. With regard to imports, although China has a large area of arable land, it has a large population and a small per capita area of arable land, which causes China having to rely on imports to meet domestic demands.
Compared to maize and rice, the soybean and wheat trade networks exhibit sparser connectivity and lower density (Fig. 1f), manifested by comparable numbers of participating economies but fewer trade relationships (Fig. 1a,b). For both of them, the first quartile is zero in the logarithmic-transformed boxplots of $k^{\mathrm{out}}$ and $s^{\mathrm{out}}$ (Fig. 2f,h), indicating their more concentrated exporting characteristics. The difference, however, is that trade volume of wheat varies less among importers and is more evenly distributed (Fig. 2g). Moreover, the soybean trade network exhibits substantial heterogeneity in the size of participating economies, indicating highly monopolistic characteristics by a few economies such as the United States, Brazil, and China (Fig. 2c).
@@
Furthermore, we calculate efficiency, redundancy, $\alpha$ values, and resilience of the international food trade networks from 1986 to 2022 (Fig. 3a-d). Overall, trade networks of maize, rice, and wheat are more resilient. Separately, maize and wheat exhibit parallel enhancements in both efficiency and redundancy, resulting in the better trade-off within the $\alpha$ framework. Rice exhibits enhancement in redundancy but fluctuation in efficiency, resulting in a slowly declining resilience in recent years. Conversely, the soybean trade network is not only more vulnerable, its resilience fluctuates more significantly. As is known, approximately 87% of soybean export quantities originated from Brazil and the United States in 2022. Due to the excessively high concentration of exporters, most economies, except for the main exporters, remain relatively isolated, with limited alternative trade pathways. Consequently, the soybean trade network lacks the capacity for development and both efficiency and redundancy, and exhibits an insufficient trade-off within the $\alpha$ framework, ultimately resulting in lower resilience. Notably, $\alpha$ values of all networks are less than $1/{\mathrm{e}}$, resulting in the highly positive correlation between efficiency and resilience (Fig. 3e) and the negative correlation between redundancy and resilience (Fig. 3f). It indicates that the current trade networks possess sufficient redundancy but require enhanced efficiency, especially for soybean.
4 Structural drivers of resilience of the international food trade network
4.1 Correlations between topological indicators and resilience
We investigate the correlations between topological indicators of the international food trade networks and their resilience and also calculate the correlation coefficients between topological indicators and both efficiency and redundancy, as shown in Fig. 4.
@@ 4.1 Correlations between topological indicators and resilience
Figure 4: a-i, Scattered distribution of resilience and topological indicators. In this figure, the red dotted line indicates $1/{\mathrm{e}}$. j, Kendall’s correlation coefficients between resilience and topological indicators, between efficiency and topological indicators, and between redundancy and topological indicators. In this figure, $R_{m}$, $R_{r}$, $R_{s}$, and $R_{w}$ denote resilience of trade networks of maize, rice, soybean, and wheat, respectively. $e_{m}$, $e_{r}$, $e_{s}$, and $e_{w}$ denote efficiency of trade networks of maize, rice, soybean, and wheat, respectively. $_{m}, rr$$_{r}$, $r_{s}$, and $r_{w}$ denote redundancy of trade networks of maize, rice, soybean, and wheat, respectively.
Across all staple foods, we find that modularity exhibits a stronger positive correlation with both efficiency and resilience due to the current networks being more redundant (more details are provided in Appendix B). This findings can also explain why more resilient states of international trade are associated with stronger community structures (Garlaschelli and Loffredo, 2005; Li et al., 2024). For maize and wheat, for example, the correlation coefficients between resilience and modularity are 0.72 and 0.80, while those between efficiency and modularity are 0.71 and 0.68, respectively (Fig. 4j). A higher modularity reflects denser intra-community flows relative to inter-community flows, corresponding to processes of regional integration that lead to stronger inter-dependence among economies within a community and, in turn, enhanced overall network efficiency (de Raymond et al., 2021). Within our framework, the resilience of the current networks is positively correlated with their efficiency, so higher modularity implies higher resilience.
@@
The correlation between resilience and other topological indicators varies across different foods. For example, the clustering characteristics are positively correlated with resilience for maize, soybean, and wheat. According to Fig. 4j, in trade networks such as those for maize and wheat, evident clustering characteristics enhance trade efficiency (e.g., through regional alliances and international cooperation organizations, such as NAFTA [^6] NAFTA, The North American Free Trade Area. and MERCOSUR [^7] MERCOSUR, The Southern Common Market., and visualizations are provided in Appendix C). Together with modularity, these features jointly contribute to sustaining network resilience. However, resilience of the rice trade network appears to decline with increases in both clustering coefficients and the network density, potentially due to its inherent network characteristics and the nature of rice as a commodity. On one hand, the rice network is heavily concentrated around a few major exporters (such as India, Thailand, and Vietnam) and importers (such as China and the Philippines), forming a single-core structure of Asia. Despite denser connectivity, this over-centralized architecture has exacerbated network vulnerability. Meanwhile, while geographic proximity suggests potential trade alliances, the lack of strategic complementarity, exemplified by competitive underpricing between Vietnam and Thailand, undermines cluster resilience by fostering internal fragmentation. Additionally, as an inelastic staple commodity with low substitution elasticity, the rice trade network faces catastrophic systemic collapse once major Asian producers implement export restrictions to ensure domestic food security (e.g., India’s 2022 export ban for rice).